OpenStudy (anonymous):

What are the minimum,first quartile,median,third quartile,and the maximum of the data set?55,62,61,54,68,72,59,61,70 pls help thnx...

OpenStudy (jack1):

median is the middle result, if u had 5 results in the data: 1, 79, 15, 78, 16 rearrange from highest to lowest: 1, 15, 16, 78, 79 median is the middle one, so 16 in this case even thought the mean is more like 38 ish

OpenStudy (jack1):

u following so far?

OpenStudy (jack1):

@Lisad9708

OpenStudy (anonymous):

I'm trying I'm kinda lost I have no clue just keep coming ill try my best to understand

OpenStudy (jack1):

ok, so here's your data: 55,62,61,54,68,72,59,61,70 and here's your data, re-arranged from highest to lowest: 54 55 59 61 61 62 68 70 72 so which is the middle number in this?

OpenStudy (anonymous):

61?

OpenStudy (jack1):

yep, so there's 61, then there's 4 numbers on either side The median is the middle value of the data set. The lower quartile (Q1) is the median of the lower half of the data set.

OpenStudy (jack1):

so, median equation: median term = half of (n+1) so our data set has 9 values median term = half of (9+1) = half of 10 = 5th term in the data set (when arranged from highest to lowest) you guessed it right before it was 61 1st quartile term = quarter of (n+1) so our data set has 9 values 1st quartile term = 1/4 of (9+1) = 1/4 of 10 = 2.5th term = average of 2nd and 3rd values

OpenStudy (jack1):

so what does that equal?

OpenStudy (jack1):

or at least: what are the 2nd and 3rd values...?

OpenStudy (anonymous):

Do I have square root or divide or times it

OpenStudy (anonymous):

Third value is 68?

OpenStudy (jack1):

no, none of the above yet here's your data, arranged from highest to lowest: 54 55 59 61 61 62 68 70 72 what are the 2nd and 3rd values?

OpenStudy (anonymous):

55,59?

OpenStudy (jack1):

yep, perfect now whats the average of them? (mean average)

OpenStudy (jack1):

= (55+59)/2 \[mean = \frac{(55 + 59)}{ 2 }\]

OpenStudy (anonymous):

When u said average do I add them 2gether?

OpenStudy (jack1):

yes, then divide by 2

OpenStudy (anonymous):

Pls don't give up on me I'm kinda slow its been awhile n math is not my thing.

OpenStudy (jack1):

that's ok, u should still be able to add 2 numbers together though and then divide that answer by 2

OpenStudy (jack1):

" n math is not my thing. " dude i was the same, but then i stopped thinking of it as numbers and more as logic puzzles, and it seemed to get easier

OpenStudy (anonymous):

57

OpenStudy (jack1):

so 55+59 = 114 114/2 = 57 so your first quartile is 57

OpenStudy (jack1):

spot on dude

OpenStudy (jack1):

now: 3rd quartile term = 3 quarters of (n+1) so our data set has 9 values 3rd quartile term = 3/4 of (9+1) = 3/4 of 10 = 7.5th term = average of 7th and 8th values

OpenStudy (jack1):

so from highest to lowest: 54 55 59 61 61 62 68 70 72 what are the 7th and 8th values?

OpenStudy (anonymous):

68+70?

OpenStudy (jack1):

yep now divide that by 2

OpenStudy (anonymous):

69

OpenStudy (jack1):

yep, that's your 3rd quartile

OpenStudy (jack1):

now, ur max and min are your highest and lowest numbers in the data set 54 55 59 61 61 62 68 70 72 shoot

OpenStudy (anonymous):

72

OpenStudy (jack1):

that's ur max done, props

OpenStudy (anonymous):

72-54

OpenStudy (jack1):

yeah, and that's pretty much it for that q, yeah?

OpenStudy (anonymous):

So minimum is 18 first quartile is 57 third quartile is 69 maximum is 72 I didn't get the medians

OpenStudy (jack1):

median was 61, first quartile 57, 3rd quartile 69 max 72... where did 18 come from...?

OpenStudy (anonymous):

Oh I thought there is a minimum

OpenStudy (jack1):

minimum = lowest number in data set = 54

OpenStudy (jack1):

ah, i getcha now, range = 18

OpenStudy (jack1):

also, don't forget to close off your questions afterwards. button beside the "bump" at the top

OpenStudy (anonymous):

Thnx for ur help I have some more question if its ok with u thnx again

OpenStudy (jack1):

sure, this ones getting a bit long, close and start a new one...?